at white plains bowl the charge per game is $2.50 and it costs $2.00 for shoe rental. at homefield bowl the cost is $2.00 per game and the shoe rental is $4.00. they are wondering for how many games will the cost to bowl be the same at both places. use elimination to solve.
Answer by mananth(16949)

Let the number of games played be x
White plain bowl
2+2.50x is the cost
Home field
4 + 2x
..
equate them to find the games
2+2.50x = 4+2x
-2x
2+2.50x-2x=4
-2
2.50x-2x=4-2
0.5x=2
/0.5
x=2/0.5
x= 4 games their costs are equal
...
m.ananth@hotmail.ca